(I just need part B)

Consider the reaction

A+2B⇌C

whose rate at 25 ∘C was measured using three different sets of initial concentrations as listed in the following table:

Trial [A]
(M)
[B]
(M)
Rate
(M/s)
1 0.50 0.030 9.0×10−3
2 0.50 0.060 1.8×10−2
3 1.00 0.030 3.6×10−2

Part A

Part complete

What is the rate law for this reaction?

Express the rate law symbolically in terms of k, [A], and [B].

rate = k[A]2[B]

Part B

Calculate the initial rate for the formation of C at 25 ∘C, if [A]=0.50M and [B]=0.075M.

Express your answer to two significant figures and include the appropriate units.
Answer #1

Part B :

Rate of reaction = k.[A]2[B]

9.0×10-3M.s-1 = k ×(0.50 M)2(0.50 M)

k = rate constant = 7.2×10-2 M-2.s -1

Initial rate of reaction = k.[A]2[B] = 7.2×10-2 M-2.s-1 × (0.50 M)2.(0.075 M) = 0.135×10-2 M/s

Initial rate of reaction = initial rate for formation of C = 1.35×10-3 M/s

= 1.4×10-3 M/s. (Answer)